### v1

# Data-driven optimization of pixelated CdZnTe spectrometers for uranium enrichment assay

Abstract
In recent work [Vavrek et al. (2025)], we developed the performance optimization framework spectre-ml for gamma spectrometers with variable performance across many readout channels. The framework uses non-negative matrix factorization (NMF) and clustering to learn groups of similarly-performing channels and sweep through various learned channel combinations to optimize the performance tradeoff of including worse-performing channels for better total efficiency. In this work, we integrate the pyGEM uranium enrichment assay code with our spectre-ml framework, and show that the U-235 enrichment relative uncertainty can be directly used as an optimization target. We find that this optimization reduces relative uncertainties after a $30$-minute measurement by an average of $20\%$, as tested on six different H3D M400 CdZnTe spectrometers, which can significantly improve uranium non-destructive assay measurement times in nuclear safeguards contexts. Additionally, this work demonstrates that the spectre-ml optimization framework can accommodate arbitrary end-user spectroscopic analysis code and performance metrics, enabling future optimizations for complex Pu spectra.
I Introduction
The H3D M400 gamma spectrometer [1] is being adopted by the International Atomic Energy Agency (IAEA) as its primary in-field uranium enrichment non-destructive assay (NDA) technology [2, 3, 4]. The M400 features four CdZnTe (CZT) crystals each pixelated to an $11\times 11$ grid, and offers medium resolution gamma spectroscopy at room temperature in a compact form factor. In previous work, we showed that the spatial variations in detector performance within the CZT crystal volumes could be exploited to optimize spectroscopic performance metrics [5]. In particular, a balance can be found between identifying and rejecting poorly-performing detector regions and the associated loss of efficiency. Because of the large number of possible voxel combinations (${\sim}10^{7285}$), computing the globally optimum set of voxels to include is infeasible, and approximate, data-driven methods were developed. Several of the example optimizations therein focused on minimizing the relative uncertainty in a peak fit amplitude parameter as proofs of concept in lieu of more advanced spectral performance metrics.
In this work, we have integrated our spectroscopic optimization software spectre-ml with GEM [6], the actual software used by the IAEA to conduct uranium NDA measurements. We show that the relative uncertainty on the GEM enrichment calculation can directly be used as a spectre-ml optimization target, i.e., that spectre-ml can improve the statistical confidence of the IAEA’s in-field enrichment measurements without increasing measurement times.
This paper is structured as follows: Section II covers the integration of the spectre-ml and GEM codes at a high level, discussing recent improvements to the former to better enable arbitrary end-user workflows, and providing an overview of the GEM workflow itself. It also describes the uranium measurements and optimization parameter sweeps used for analysis. Section III gives the results of six GEM-based spectre-ml optimizations for six M400 units, showing that the enrichment relative uncertainty can be directly used as an optimization target. Section IV then provides additional interpretation of the results, discusses some limitations, and suggests avenues for further work. Finally, the Appendix gives a pseudocode overview of how an end-user can integrate their own spectrum analyses into the spectre-ml optimization framework.
II Methods
II-A Integration of the spectre-ml and pyGEM codes
The spectre-ml [5, 7] code is a Python package for optimizing the performance of many-channel gamma spectrometers by intelligently rejecting data from poorly-performing detector channels (i.e., spatial regions). It uses non-negative matrix factorization (NMF) [8, 9] and unsupervised clustering algorithms to learn groups of voxels with similar performance, compute performance metrics across various voxel cluster combinations, and then sweep over hyperparameters such as clustering algorithm, number of clusters, number of NMF components, etc., to find the best overall set of voxels to include. Previous versions of spectre-ml were restricted to single Doniach [10] or Gaussian peak fit workflows using the becquerel library for peak fits. In this work we have generalized the spectre-ml software to abstract out the dependence on becquerel and enable arbitrary user-defined spectrum analyses, in particular those based on GEM. The pyGEM [11] code is a Python interface to the General Enrichment Measurements (GEM) code [6] used by the IAEA for uranium enrichment NDA. It fits measured spectra in the ${\sim}120$$270$ keV ROI using a set of known U-235 emission lines, forward scatter, and high-energy downscatter profiles, and has recently been updated to handle the asymmetric peak shapes of CZT. In contrast to our Doniach fit becquerel workflow, pyGEM uses a triple-Gaussian model for each peak in CZT—one primary peak, one low-energy tail, and one high-energy tail. The user provides pyGEM a spectrum from a sample of known U-235 enrichment, and it will then compute a linear calibration between the fit net count rate in the $185.7$ keV peak and the U-235 atom abundance that can be used to compute the enrichment from an unknown sample spectrum. It is also possible to incorporate correction factors for certain changes between the calibration and sample spectra such as changes in the sample matrix (e.g., U metal vs. U3O8), Al and steel wall thicknesses, etc.
To integrate the two codes, some spectre-ml updates were required to handle both the becquerel and pyGEM workflows. As shown in the pseudocode in the Appendix, spectre-ml now provides the abstract base classes SpectrumAnalyzer and RankingMetric. The user must provide concrete subclasses implementing SpectrumAnalyzer.analyze and RankingMetric.calc, where the latter method takes a spectrum analyzed (e.g., fit) by the former and extracts a single float metric such as the relative uncertainty in a fit parameter (becquerel) or in an enrichment (pyGEM) calculation. The usual spectre-ml analysis then proceeds and the various voxel selections tested are then ranked by these metric values, with the best metric value indicating which voxel selection to use for the given application defined by the input spectra and choice of spectrum analysis and ranking metric. We note that no pyGEM code changes were required, indicating that spectre-ml end-users can relatively easily slot in their own analysis routines.
II-B Uranium enrichment assay optimization
Multiple U3O8 standards measurements were performed with six US National Laboratory M400 detectors (Brookhaven, Idaho, Los Alamos, Oak Ridge, Pacific Northwest, and Sandia National Laboratories) [12] and provided to Lawrence Berkeley National Laboratory for analysis with spectre-ml. In this work we use the $30$-minute collimated M400 measurements of Standard Reference Material (SRM) 969 [13] samples with label U-235 enrichments of $1.94$ wt% and $4.46$ wt% (certified values of $(1.9420\pm 0.0014)$ wt% and $(4.4623\pm 0.0032)$ wt%). The latter is used as the known-enrichment sample for the pyGEM activity calibration while the former is used as the unknown sample to be assayed.
The relative uncertainty in the U-235 enrichment is a crucial performance metric for U-235 NDA as it is directly used in computing the operator-inspector difference (OID) in terms of $z$-scores or “sigmas”, and used to compare uncertainties against the International Target Values (ITVs) for the GEM-based verification method [14]. A lower relative uncertainty on the measured U-235 enrichment increases the statistical power of the measurement to detect diversions from declared enrichments and thus potential treaty violations. Since the statistical component of the relative uncertainty decreases with the square root of measurement time, long measurement times may be required for high statistics; longer measurement times are however not always an option for inspectors performing verifications in the field. Optimizing for lower relative uncertainty can therefore improve the time-efficiency of IAEA NDA tasks.
In the optimization examples of Section III, the spectre-ml optimization considers four different classes of algorithms for selecting voxel cluster combinations: machine learning (ML), heuristic, greedy, and random voxel selections. The ML parameter sweeps consist of $n_{\text{clus}}=2$66 clusters (Gaussian Mixture only), $n_{\text{comp}}=2$66 NMF components, and $\alpha_{W}=0.0$ (i.e., no NMF regularization). The heuristic algorithms are the equal-depth and edge-and-anode algorithms introduced in Ref. [5], with $n_{\text{clus}}=2$66 and $n_{\text{anode}}=15$, $n_{\text{edge}}=1$, respectively. The greedy detector algorithm is used, but not the greedy pixel, depth bin, or voxel algorithms due to memory constraints, and a single sample is run for each of the random pixel, depth bin, and voxel clusterers with $n_{\text{clus}}=2$66.
III Results
Fig. 1 shows the results of the spectre-ml + pyGEM U-235 enrichment relative uncertainty optimization using the PNNL M400 uranium standards dataset. The enrichment relative uncertainty of $1.06\%$ from the bulk (unoptimized) spectrum is improved to $0.93\%$ (a relative improvement of $12\%$) when using $n_{\text{comp}}=3$ and retaining $1$ out of $3$ Gaussian Mixture clusters. This voxel mask reduces the $185.7$ keV net peak area from $8.22\times 10^{5}$ cps to $4.45\times 10^{5}$ cps (an efficiency reduction to $54\%$ of the unoptimized detector) and slightly degrades the peak full width at half maximum (FWHM) from $1.18\%$ to $1.20\%$, but improves the overall fit quality from $2.70$ to $1.93$. The top $5$ masks achieve similar final metrics of $0.93\%$$0.94\%$. Several of the clustering algorithms (notably the edge-and-anode) clusterer never achieve results better than the bulk value of $1.06\%$.
Figure 1: Uranium enrichment relative uncertainty optimization example with the PNNL M400 detector. Top left: best cluster labels. Top right: best cluster mask. Center left: top 55 spectra ranked by enrichment relative uncertainty. Center right: best spectrum from each class of clustering algorithm. Bottom left: pyGEM fit to the bulk spectrum. Bottom right: pyGEM fit to the best spectrum.
Similar or better spectre-ml + pyGEM improvements in the U-235 enrichment relative uncertainty are observed across all six M400 detectors tested—see Table I. The metric relative improvements range from $12\%$ to $26\%$, with a mean of $20\%$. Five out of six detectors are best improved by a GaussianMixture clusterer with $2$ out of $3$ clusters retained, while the PNNL detector removes one additional cluster and the BNL detector sees best results from the random depth bin clusterer—see the best cluster masks in Fig. 2. The relative efficiencies after voxel cluster removal range from $54\%$ (PNNL) to $70\%$ (BNL), with the remaining four detectors more tightly grouped between $61\%$ and $68\%$. All optimizations produce very small FWHM degradations, on average changing from $1.15\%$ (bulk) to $1.21\%$ (best), for reasons that are not yet known. Finally, the pyGEM fit quality improves in all six optimizations, from an average of $2.81$ to $1.71$.
detector BNL INL LANL ORNL PNNL SNL mean
best model ($n_{\text{comp}};n_{\text{clus}}$) $6;2\text{ of }3$ $5;2\text{ of }3$ 2;2 $2\text{ of }3$ $4;1\text{ of }3$ $4;2\text{ of }3$
186 keV rel. eff. $70\%$ $62\%$ $61\%$ $68\%$ $54\%$ $65\%$ $63\%$
U-235 wt%, bulk $1.974\pm 0.020$ $1.935\pm 0.020$ $1.951\pm 0.020$ $1.938\pm 0.019$ $1.956\pm 0.021$ $1.875\pm 0.021$ $1.938\pm 0.008$
U-235 wt%, best $1.959\pm 0.018$ $1.937\pm 0.015$ $2.004\pm 0.016$ $1.950\pm 0.014$ $1.962\pm 0.018$ $1.901\pm 0.016$ $1.952\pm 0.007$
metric improvement $13\%$ $24\%$ $21\%$ $24\%$ $12\%$ $26\%$ $20\%$
186 keV FHWM, bulk $12\% 1.1$$3\% 1.1$$15\%$ $20\% 1.1$$8\% 1.1$$0\% 1.1$$15\%$
186 keV FHWM, best $1.14\%$ $1.21\%$ $1.24\%$ $1.32\%$ $20\% 1.1$$14\%$ $1.21\%$
pyGEM fit quality, bulk $2.78$ $2.78$ $2.73$ $2.47$ $2.70$ $3.37$ $2.81$
pyGEM fit quality, best $2.05$ $1.55$ $1.62$ $1.37$ $1.93$ $1.76$ $1.71$
TABLE I: Summary of spectre-ml + pyGEM U-235 enrichment assay optimization results across six M400 detectors.
Figure 2: Best cluster mask from each detector optimization.
The analyses above were run with the full $30$-minute dwell time datasets, where systematic fit uncertainty is expected to dominate over statistical counting uncertainty. To expand this analysis, we also test the performance of these long-dwell-computed masks on shorter sub-samples of the same data. Fig. 3 shows the U-235 enrichment (wt%) and its relative uncertainty (the direct optimization metric) as a function of dwell time $t$ for three of the six detectors. As expected, the relative uncertainty decreases with increasing dwell time, for both the bulk and optimized results. The initial optimized relative uncertainty tends to be larger than the initial bulk value, presumably due to worse counting statistics from the lower efficiency, but the rate of decrease in the optimized curves is faster, leading to a crossover time, $t^{*}$, where the optimized relative uncertainty first falls below the bulk result. As shown in Table II, these crossover times range from $17$ s to $117$ s, with an average of $49$ s, depending on the detector. Similarly, we can quantify the speedup factor, $f_{t}$, that could be achieved by dwelling with the optimized model only until the relative uncertainty from the bulk dataset at time $t$ is reached. For the full $30$-minute datasets, these speedups $f_{30}$ range from $5.2\times$ to $15.4\times$, with an average of $10.3\times$, and are also given in Table II. For a shorter $5$-minute dwell time, the speedups $f_{5}$ range from $2.3\times$ to $5.2\times$, with an average of $3.8\times$, indicating that the optimized models can achieve the same relative uncertainty as a $5$-minute bulk measurement in about a quarter of the time.
Figure 3: Trends in pyGEM analysis metrics vs dwell time $t$, using the optimum long-dwell-computed mask from the INL (top row), SNL (middle row), and LANL (bottom row) detectors. Left column: U-235 enrichment (wt%). The optimized (orange) points are slightly offset along the $x$-axis from their corresponding bulk (blue) points for visual clarity. The gray band shows the SRM 969 certification of $1.9420\pm 0.0014$ wt% [13]. Right column: U-235 relative uncertainty (%).
detector crossover time $t^{*}$ [s] speedup factor $f_{30}$ $f_{5}$
BNL 23 8.9 3.0
INL 39 11.7 5.2
LANL 68 8.9 3.0
ORNL 30 11.7 3.9
PNNL 117 5.2 2.3
SNL 17 15.4 5.2
mean 49 10.3 3.8
TABLE II: Summary of time efficiency improvements
In addition to the improvement in the relative uncertainties in Fig. 3—the precisions of the enrichment calculations—it is also important to discuss the effects on the accuracies. The INL, SNL, and LANL detector datasets in Fig. 3 were specifically chosen to highlight three different cases: the bulk detector accuracy is essentially unchanged by the optimization (INL); it is slightly improved by the optimization (SNL); and it is slightly degraded by the optimization (LANL). Similarly, Fig. 4 summarizes the U-235 enrichment assay results for all six detectors. The LANL M400 result appears to be an outlier that pulls the six-detector mean above the sample’s declared enrichment band, but other than the LANL detector, each detector result agrees within error bars. Three of the six detectors have computed U-235 enrichments closer to the declared value without using the spectre-ml optimization, and three of six are closer with it. With or without the spectre-ml optimization, the computed enrichments can fall on either side of the declared value—but with the optimization, the uncertainties are improved without any obvious bias.
Figure 4: Summary of U-235 enrichment assay results with (orange squares) and without (blue circles) the spectre-ml + pyGEM optimization.
IV Discussion
The results of Section III retain a handful of limitations and present opportunities for future work. First, the spectre-ml parameter sweeps are much smaller than those used in Ref. [5], using only 2266 clusters and NMF components, only the Gaussian Mixture clusterer (among the ML options), and no NMF regularization. It is likely that the observed $20\%$ average relative improvement in the U-235 enrichment relative uncertainty metric would improve if the parameter sweeps were expanded. Currently, we are compute-limited by the worse performance of the underlying scikit-learn libraries on the Windows machine required for the pyGEM code, with even these smaller parameter sweeps taking >1{>}1 hr wall time on a $4$-core 11th Gen Intel Core i7-1185G7 processor. Performance is also degraded by the need to run both NMF and fitting routines on spectra of $O(1000)$ bins for the entire U-235 region rather than the $O(100)$ bins for a single photopeak. Better parallelization, either in the scikit-learn routines or over individual parameter combinations, could greatly improve the speed of the optimization.
We also note that the observed $20\%$ average relative improvement is smaller than the ${\sim}3\times$ improvement observed in the U-235 peak analysis of Ref. [5], as expected, due to the much lower systematic fit error arising from the more complex spectrum fit (which includes the $195$ keV peak) in pyGEM. This re-emphasizes the importance of carefully defining the performance metric to be used to avoid specification gaming.
The similarity of four out of six best voxel removal models in Table I and therefore masks in Fig. 2 suggests that a single common mask could be applied to improve many M400 detectors without having to specifically optimize each detector, even despite inter-detector performance differences. It would be valuable in the future to quantify this level of mask transferability among detectors.
Our analysis here has been limited to using the $4.46\%$-enriched U-235 standard as a calibration standard for assaying the $1.94\%$-enriched sample. In the future, it could be interesting to repeat the analysis for the reverse combination, and to include the $20.11\%$-enriched measurements from the same collection of datasets. Similarly, other safeguards measurement scenarios such as UF6 cylinders could be considered.
Finally, we note that the ease with which pyGEM was integrated with spectre-ml bodes extremely well for future integrations of other analysis codes. For instance, FRAM [15] could be integrated to test the ability of spectre-ml to improve the challenging analysis of multiple densely-packed photopeaks in Pu spectra.
V Conclusions
We have integrated the pyGEM uranium enrichment analysis code with the spectre-ml spectroscopic optimization framework, and shown that the uranium enrichment relative uncertainty can be directly used as an optimization target. This spectre-ml + GEM workflow provides a $20\%$ relative improvement (averaged over six H3D M400 detectors) in the U-235 enrichment relative uncertainty in $30$-minute long measurements, and retains similar enrichment accuracy to the unoptimized pyGEM-only results. These reductions in enrichment relative uncertainty can lead to significant time savings, with the optimized results achieving the same overall relative uncertainty as the unoptimized $30$-minute measurements around $10\times$ faster.
The following pseudocode provides a basic outline of the application programming interface (API) that an end-user would interact with to integrate their own analysis code with the spectre-ml framework.
# spectre-ml framework: abstract classes
class RankingMetric(ABC):
class SpectrumAnalyzer(ABC):
# user code: concrete becquerel workflow
class BecquerelSpectrumAnalyzer(SpectrumAnalyzer):
def analyze(self):
compute_single_peak_fits()
class BecquerelRankingMetric(RankingMetric):
def calc(self) -> float:
return fit_param_rel_unc
# user code: two concrete pyGEM workflows
class GEMSpectrumAnalyzer(SpectrumAnalyzer):
def analyze(self):
compute_activity_calibration()
compute_U235_region_fits()
class GEMEnrichmentMetric(RankingMetric):
def calc(self) -> float:
return enrichment_rel_unc
class GEMFittingMetric(RankingMetric):
def calc(self) -> float:
return fit_chi_squared
# user code: example GEM analysis
ms = spectre_ml.ModelSelector(
spectrum_analyzer=GEMSpectrumAnalyzer(),
ranking_metric=GEMEnrichmentMetric(),
voxel_spectra=voxel_spectra,
n_clusters=[2, 3],
n_nmf_components=[2, 3],
ms.fit_models()
results = ms.evaluate_models()
results.plot()
Acknowledgments
The work presented in this paper was funded by the National Nuclear Security Administration of the Department of Energy, Office of International Nuclear Safeguards. This work was performed under the auspices of the U.S. Department of Energy by Lawrence Berkeley National Laboratory (LBNL) under Contract DE-AC02-05CH11231. The U.S. Government retains, and the publisher, by accepting the article for publication, acknowledges, that the U.S. Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this manuscript, or allow others to do so, for U.S. Government purposes.